OpenStudy (anonymous):

What is the equation of a line that passes through (5, 3) and (-5, 3)?

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

@completeidiot i used the slope formula but i didn't get the right answer?

OpenStudy (anonymous):

ok, what did you get as a slope?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

ok, now take the slope, and a single point plug those into the point slope equation and show me what you get

OpenStudy (anonymous):

using (5,3) so y-3=0(x-5)

OpenStudy (anonymous):

@completeidiot

OpenStudy (anonymous):

ok now distribute the zer